suppose a uniform spherical star of mass M and radius R collapses to a uniform sphere of radius 10^-5 .If the original star had a rotation rate of 1 rev each 25 days(as does the sun) ,what will be the rotation rate of the resulting object
1
Expert's answer
2020-04-21T19:19:20-0400

As per the question,

Mass of the spherical star=M

Radius of the spherical star= R

Radius of the other star"=10^{-5}R"

rotational speed of the first star"\omega_1= 1/25 rev/day"

rotational speed of the second star "\omega_2=?"

Now, applying the conservation of angular momentum,

"I_1\omega_1=I_2\omega_2"

"\Rightarrow \omega_2=\dfrac{I_1\omega_1}{I_2}=\dfrac{10^{10}}{25}=4\times 10^8 rev/day"

"\omega_2=4.6\times 10^3 rev/sec"